# Investigating symptom duration using current status data: a case study of post-acute COVID-19 syndrome

### v2

@@ Acknowledgements
Figure S4: Confidence interval coverage using the CIR procedure for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from smallest to largest degree of nonresponse. The top and bottom rows denote the standard CIR method using only complete cases and the extended CIR method accounting for nonresponse. The $x$-axis corresponds to sample size. The black line denotes the nominal 95% coverage level.
S3.2   Cox proportional hazards regressions simulations
@@
Data were generated as in the previous set of experiments. To evaluate the performance of the bootstrap, we fit a correctly specified Cox model (i.e., an additive model with only first-order terms) on each of $B$ bootstrap resamples of the dataset. We constructed confidence intervals for each regression coefficient using either a normal approximation with standard deviation equal to the standard deviation of the point estimates across the $B$ replicates or a percentile bootstrap. We repeated this procedure 1000 times and calculated the empirical confidence interval coverage for each method of interval construction. We generated datasets of size $n=\{500,1000,1500,2000\}$ and used bootstrap replicates $B=\{100,250,500,1000\}$. The empirical coverage for the coefficients corresponding to $W_{1},W_{2}$ and $W_{3}$ in each of these settings is displayed in Figures S5S7.
From Figures S5S7, we observe that, under all nonresponse mechanisms, the normal approximation bootstrap performs well, with empirical coverage within Monte Carlo error of the nominal level of 0.95 across both sample sizes and number of bootstrap replicates. The percentile bootstrap produces somewhat anticonservative confidence intervals, although its performance tends to improve with a larger number of bootstrap replicates.
@@
S4.   Supplementary information for HCT data analysis
@@
S4.1   Eligibility criteria for HCT enrollment
@@
S4.2   Algorithm library for Super Learner
For data analyses and simulations Super Learner [33] was implemented with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines [48], generalized additive models [49], random forests [50], and gradient boosted trees [51].
S4.3   Supplementary figures
Figure S8 shows the distribution of survey response time $Y$ among HCT survey respondents.

### v3

@@ Acknowledgements
Figure S4: Confidence interval coverage using the CIR procedure for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from smallest to largest degree of nonresponse. The three lines in color denote the standard CIR method using only complete cases, the extended CIR method accounting for nonresponse, and the NPMLE. The $x$-axis corresponds to sample size. The black line denotes the nominal 95% coverage level.
S4.2   Robustness and stability simulations
In this section, we describe a simulation study performed to evaluate the robustness and stability of the extended CIR procedure. Data were generated as described above, under Scenario 3, with $c_{0}=1.65$. We generated datasets of size $n\in\{500,1000,2500,5000\}$.
We used the extended CIR procedure, with several different combinations of nuisance estimators, to estimate the survival function of $T$. The conditional density $\pi$ was estimated using either (1) a misspecified parametric log-normal accelerated failure time model including only linear terms, or (2) haldensify, an implementation of a nonparametric approach based on the highly adaptive lasso (Hejazi et al.,, 2022). The outcome regression nuisance function $\mu$ was estimated using either (1) logistic regression, (2) GAMs (Hastie,, 2017), (3) gradient boosted trees (Chen and Guestrin,, 2016), or (4) Super Learner (van der Laan et al.,, 2007) with a library as described in the previous experiment. For comparison, we also included the true form for $g$ and $\mu$ (i.e., in lieu of estimating the nuisance parameters, we plugged in their known values). As long as either of the two nuisance estimators is sufficiently flexible to consistently estimate its target, we expect the extended CIR procedure to yield a consistent estimator. Furthermore, if the nuisance estimation rate is fast enough, as described in Section S1.2, we expect the resulting extended CIR estimator to not only be consistent, but to have a large-sample distribution that does not depend on the particular nuisance estimator employed. This large-sample distribution be identical as when the nuisance parameters are set to their true values. For each sample size, we conducted 500 simulations replicates; all 15 possible combinations of nuisance estimators were applied to each simulated dataset. We examined pointwise empirical bias and pointwise empirical variance. Furthermore, we examined the absolute difference between survival function estimates computed on each dataset, comparing (1) using known nuisance parameters versus using the haldensify–Super Learner combination, and (2) using the haldensify–Super Learner combination versus using the haldensify–gradient boosted trees combination. These two comparisons quantify the sensitivity of the proposed procedure to the choice of nuisance estimator when a flexible method is used in place of the ‘ideal’ method using known nuisance parameter values and when two distinct flexible methods are used, respectively.
From Figure S5, we observe that when $\pi$ is estimated using a misspecified parametric model and $\mu$ is estimated using either GLM or GAM, the extended CIR estimate has bias that does not shrink with increasing sample size. In these scenarios, neither of the two nuisance estimators can consistently estimate its target — the parametric density model and GLM are both misspecified, and the underlying outcome regression is not additive — yielding an overall CIR estimator that appears inconsistent as well. In all other cases, the overall bias is generally small. Using gradient boosted trees or Super Learner to estimate $\mu$ and haldensify to estimate $\pi$ results in similar performance to using the true, known values for the nuisance parameters. From Figure S6, we observe that the empirical variance of the CIR estimator exhibits similar patterns regardless of the nuisance estimators used.
Figure S5: Pointwise bias of extended CIR procedure using various nuisance estimators. The columns represent different estimators for the conditional density $\pi$, and the rows represent different estimators for the outcome regression $\mu$. The $x$-axis displays the true value of the distribution function. The black line denotes zero bias.
Figure S6: Pointwise variance of extended CIR procedure using various nuisance estimators. The columns represent different estimators for the conditional density $\pi$, and the rows represent different estimators for the outcome regression $\mu$. The $x$-axis displays the true value of the distribution function.
In Figure S7, we plot the median, 5th percentile, and 95th percentile absolute difference in survival function point estimates computed using haldensify and either gradient boosted trees or Super Learner, with the quantiles calculated across the 500 simulation replicates. This quantifies the stability of the CIR procedure; gradient boosting and Super Learner will produce different survival function estimates even when applied to the same data, but smaller differences indicate greater stability and less dependence on the chosen algorithm. We see that the stability of the procedure increases with increasing sample size. At sample size 5000, the median difference is less than 0.002 across all time points. The 95th percentile absolute difference is never larger than 0.01. Even in a relatively small sample of size 500, the median absolute difference exceeds 0.01 only near the boundary of the observable support. These results suggest that, particularly in large samples, survival estimates tend to be insensitive to the choice of nuisance estimators, at least among well-performing nuisance estimators.
Figure S7: Pointwise absolute difference between estimates of the distribution function computed using either gradient-boosted trees or Super Learner for the outcome regression $\mu$, along with haldensify for the conditional density $\pi$. For each simulation replicate, each of the two nuisance estimators was applied to the simulated data and then used in the extended CIR procedure. The quantiles displayed in the plot were calculated over 500 simulation replicates.
In Figure S8, we plot the median, 5th percentile, and 95th percentile absolute difference in survival function point estimates computed using either (1) haldensify and Super Learner or (2) the true, known values of the nuisance parameters. Again, the stability of the procedure increases with increasing sample size, although in general the differences tend to be slightly larger than those between the two machine learning–based procedures in Figure S7. At sample size 500, the median difference does not exceed 0.02.
Figure S8: Pointwise absolute difference between estimates of the distribution function computed using either Super Learner and haldensify or the true, known forms for the outcome regression $\mu$ and conditional density $\pi$. For each simulation replicate, each of the two approaches was applied to the simulated data and then used in the extended CIR procedure. The quantiles displayed in the plot were calculated over 500 simulation replicates.
S4.3   Sensitivity analysis simulations
In this section, we describe a simulation study conducted to assess the performance of the proposed sensitivity analysis approach. The data were generated according to a proportional hazards model with the Frank copula used to induce dependence between $T$ and $Y_{*}$ given $W$. First, we simulated a covariate vector $W$ as in the previous experiments. Given covariate vector $W=w$, we then simulated $Y_{*}$ from a Weibull distribution with shape parameter 0.75 and scale parameter $\exp(\tfrac{2}{5}w_{1}-\tfrac{1}{5}w_{2}+\tfrac{1}{10}w_{3}).$ Then, for association parameter $\alpha$, given $W=w$ and $Y_{*}=y$, the symptom resolution time $T$ was simulated according to the Frank copula with parameter $\alpha$ and Weibull distribution with scale parameter $\exp(\tfrac{2}{5}w_{1}-\tfrac{1}{5}w_{2}+\tfrac{1}{10}w_{3})$. In this way, $T$ and $Y_{*}$ did not satisfy conditional independence, even given covariate vector $W$. We varied the association parameter $\alpha$ to yield Kendall’s $\tau$ values
$\{-0.35,-0.25,-0.15,-0.05,+0.05,+0.15,+0.25,+0.35\}\ .$
We generated datasets of size $n\in\{500,1000,2000,4000\}$. The maximum follow-up time $c_{0}$ was set to 1.65. We aimed to estimate the cumulative distribution function of $T$ over $[t_{0},t_{1}]$, with $t_{0}=0.02$ and $t_{1}=1.5$. The estimation procedure was as described in Section S4.1. We compared the performance of the extended CIR procedure using the copula-based one-step estimator $\Gamma_{P_{n},\alpha}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},\alpha,y_{0}}(Z_{i})$ to the performance of the extended CIR procedure using $\Gamma_{P_{n}}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},y_{0}}(Z_{i})$, which we expect to provide consistent estimation only under conditional independence. We refer to the latter as the ‘baseline’ extended CIR. We quantified performance in terms of empirical bias. For each simulation setting, we performed 500 simulation replicates.
From Figure S9, we observe that the copula-based estimator has substantially reduced bias compared to the baseline extended CIR estimator. The magnitude of excess bias for the baseline extended CIR estimator increases for values of Kendall’s $\tau$ further from 0.
Figure S9: Pointwise bias of extended CIR procedure using the copula-based one-step estimator versus baseline extended CIR, which requires conditional independence between symptom resolution and survey response times given measured covariates. The columns represent different values of Kendall’s rank correlation coefficient $\tau$, which describes the level of dependence between symptom resolution time and survey response time, conditional on covariates, with $\tau=0$ representing conditional independence. The $x$-axis displays the true value of the distribution function. The black line denotes zero bias.
S4.4   Cox proportional hazards regressions simulations
@@
Data were generated similarly as in the previous set of experiments; however, the scale parameter of the Weibull distribution was set to $\exp(\tfrac{2}{5}w_{1}-\tfrac{1}{5}w_{2})$, satisfying a linear proportional hazards model. To evaluate the performance of the bootstrap, we fit a correctly specified Cox model (i.e., an additive model with only first-order terms) on each of $B$ bootstrap resamples of the dataset. We constructed confidence intervals for each regression coefficient using either a normal approximation with standard deviation equal to the standard deviation of the point estimates across the $B$ replicates or a percentile bootstrap. We repeated this procedure 1000 times and calculated the empirical confidence interval coverage for each method of interval construction. We generated datasets of size $n\in\{500,1000,1500,2000\}$ and used bootstrap replicates $B\in\{100,250,500,1000\}$. For each simulation setting, we performed 500 simulation replicates. The empirical coverage for the coefficients corresponding to $W_{1},W_{2}$ and $W_{3}$ in each of these settings is displayed in Figures S10S12.
From Figures S10S12, we observe that, under all nonresponse mechanisms, the normal approximation bootstrap performs well, with empirical coverage within Monte Carlo error of the nominal level of 0.95 across both sample sizes and number of bootstrap replicates. The percentile bootstrap produces somewhat anticonservative confidence intervals, although its performance tends to improve with a larger number of bootstrap replicates.
@@
S5.   Supplementary information for HCT data analysis
@@
S5.1   Eligibility criteria for HCT enrollment
@@
S5.2   Algorithm library for Super Learner
For data analyses and simulations Super Learner (van der Laan et al.,, 2007) was implemented with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines (Friedman,, 1991), generalized additive models (Hastie,, 2017), random forests (Wright and Ziegler,, 2015), and gradient boosted trees (Chen and Guestrin,, 2016).
S5.3   Supplementary figures
Figure S13 shows the distribution of survey response time $Y$ among HCT survey respondents. Figure S14 shows the distribution of survey invitation time (time between positive test and survey invitation being sent) among all HCT survey invitees.
@@
Figure S14: Histogram of survey invitation times smaller than 120 days. Surveys were sent at least 28 days after the positive test, and 33 invitation times were larger than 120 days.